if horizontal and vertical components are equal then the angle between them is>..

Grade:9

1 Answers

Arun
6 years ago
suppose we talk about velocity (V)
then horizontal component of V =  Vx = V cos\theta
vertical component of V = Vy = V sin\theta
if they are equal then,
 V cos\theta =  V sin\theta
 tan\theta = 1
hence  \theta = 45degree